figure out what Alice's height would be multiplied by if she ate 1 ounce, 2 ounces, 3 ounces, or 4 ounces of this cake
1
Expert's answer
2020-04-01T08:25:47-0400

Alica height gets doubled on eating one ounce of this cake.

Let her height be "x."

So if she eats one ounce cake ,her height will be two times of her initial height that is "2x."

If she eats another cake,that is two ounce cake,her height will be doubled of her height before it"=2\times 2x=2^2x=4x" (four times)

If she eats another cake,that is threeounce cake,her height will be doubled of her height before it"=2\times 4x=2^3x=8x" (eight times of her initial height).

If she eats another cake,that is four ounce cake,her height will be doubled of her height before it"=2\times 8x=2^4x=16x" (sixteen times of her initial height).